# Proactive Model Adaptation Against Concept Drift for Online Time Series Forecasting

### v4

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3690624.3709210 ccs: Information systems Data stream miningccs: Computing methodologies Online learning settings
@@ 1. Introduction
Time series forecasting has been a prevalent task in numerous fields such as climate (Bi et al., 2023), energy (Boussif et al., 2023; Sun et al., 2023), retail (Böse et al., 2017), and finance (Zhao et al., 2023). Recent years have witnessed a surge of deep learning-based forecast models (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2023; Woo et al., 2024; Han et al., 2024) that take past time series observations to predict values in the next $H$ steps, where $H$ is called forecast horizon. Due to the dynamic nature of the environment, latent concepts that influence observation values (e.g., social interest (Bai et al., 2023), stock market sentiment (Li et al., 2022)) often change over time. This ubiquitous phenomenon is known as concept drift (Gama et al., 2014). In the presence of concept drift, future test data may not follow a similar data distribution as the historical training data, causing degradation in forecasting performance (You et al., 2021; Zhao et al., 2023).
@@
Online learning is commonly used to mitigate the effects of concept drift. The key consideration is to continually transform the newly observed time series into a set of training samples and adjust model parameters by minimizing the forecast errors on the new training samples. In addition to the standard fine-tuning technique, recent works (Pham et al., 2023; Zhang et al., 2023) proposed more advanced model adaptation techniques, which focus on how to effectively adapt to recent data by utilizing forecasting feedback (e.g., errors or gradients) on the new training samples. Among them, FSNet (Pham et al., 2023) monitors the gradients in previous fine-tuning, transforms them into parameter adjustments, and tailors a new forecast model to the current training samples. OneNet (Zhang et al., 2023) is an online ensembling network that generates ensemble weights to combine two forecast models and dynamically adjusts the ensemble weights and the models’ parameter weights according to the forecast errors. However, existing online learning methods overlook the fact that the ground truth of each prediction is not available immediately but is delayed after the forecast horizon.
@@
As illustrated in Fig. 1, an online learning task inherently has a $H$-step feedback delay in time series forecasting, resulting in a temporal gap (at least $H$ steps) between available training samples and the test sample. Formally, at each online time $t$, the current horizon window $\mathbf{Y}_{t}$, which has not been observed yet, has an overlap with $\mathbf{Y}_{t-H+1},\dots,\mathbf{Y}_{t-1}$. These samples cannot provide complete forecasting feedback and supervision for the performant, prevailing time series forecasting models that predict $H$-step values directly (Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024). Hence, the available training samples for online learning at time $t$ is $\mathcal{D}_{t-}=\{(\mathbf{X}_{t^{\prime}},\mathbf{Y}_{t^{\prime}})\mid t^{\prime}\leq t-H\}$, which has a temporal distance to the test sample $(\mathbf{X}_{t},\mathbf{Y}_{t})$. The temporal gap issue suggests that addressing concept drift by fitting the forecast model to the recent training samples may be insufficient, since concept drift may also occur over the $H$-step temporal gap between the recent training samples and the current test sample.
@@
To investigate the impact of this temporal gap on forecasting performance, we examine two online learning strategies using real-world time series datasets (detailed in Section 3). The first strategy fine-tunes model parameters using the latest available training sample at time $t$, which is $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$. The second strategy (Pham et al., 2023; Zhang et al., 2023) omits the feedback delay and adapts the model using $(\mathbf{X}_{t-1},\mathbf{Y}_{t-1})$, which is infeasible in practice due to future information leakage. The empirical results show that the average forecast error of the first strategy is approximately double that of the second strategy, indicating that even the latest observed time series pattern differs from the test sample notably. Moreover, the performance gap between the two strategies becomes more significant with a longer horizon, suggesting that the concept drift becomes more pronounced over longer time intervals. These findings highlight the presence of a substantial concept drift between the practical training samples and the test sample, limiting the effectiveness of current online learning techniques that adapt model parameters to potentially outdated concepts and leave the concept drift unresolved.
@@
In this paper, we aim to answer the question: how can we effectively adapt the forecast model to each test sample and boost online time series forecasting performance against ongoing concept drifts? A straightforward idea is to estimate a sketch of the latent concept of the test data and customize model parameters for the estimated concept. However, it is challenging to train a model adapter (e.g., a mapping function) that directly generates model parameters based on the estimated concept of the test sample. First, test data may bear a new concept that is out of the distribution of the historical time series. There is no experience in generating optimal parameters for such an out-of-distribution concept. Second, a parameterized model adapter that maps concepts to model parameters (or their updates) essentially lies in a parameter space of $\mathbb{R}^{m\times n}$, where $m$ is the concept embedding dimension and $n$ is the number of model parameters. Since advanced time series forecasting models often involve a high quantity of parameters (e.g., 1 million (Nie et al., 2023; Liu et al., 2024)), the model adapter is hard to optimize and easily suffers from overfitting due to the huge parameter size and limited test data.
[…]
@@
[…]
@@ 2. Definitions and Problem
Typically, in multi-step forecasting where $H>1$, there are two primary strategies to generate the predictions ${\bf\hat{Y}}_{t}$ at each time $t$. The first strategy performs iterative forecasting. That is, the model forecasts one step ahead and uses its prediction to forecast the next step, repeating this process for the entire horizon. Despite its simplicity, this strategy suffers from significant error accumulation over long horizons (Zeng et al., 2023). The second strategy adopts direct forecasting where the model makes all the predictions ${\bf\hat{Y}}_{t}$ simultaneously. Note that the two strategies require different output modules or layers in the forecast model. Recent works (Zeng et al., 2023; Nie et al., 2023) have shown that direct forecasting tends to outperform the iterative method, particularly for longer forecast horizons. Hence, this paper adopts the direct forecasting strategy where a sample $\mathbf{X}_{t}$ is considered valid for training if all the $H$-step values in $\mathbf{Y}_{t}$ are known (Chen et al., 2024).
In online forecasting scenarios, time series data are observed sequentially. Due to the dynamic nature of real-world processes, underlying data distributions are subject to constant change. Consequently, a forecast model trained on historical data may encounter difficulties when confronted with new, evolving patterns. This issue is referred to as the concept drift challenge, which can substantially affect model performance over time (Gama et al., 2014). To mitigate concept drifts, it is crucial to adapt the forecast model continuously to assimilate new concepts presented in the incoming time series. Formally, the online model adaptation problem is defined as follows.
@@
At a high level, existing model adaptation methods (Pham et al., 2023; Zhang et al., 2023; Chen et al., 2024) select some observed data $\mathcal{D}_{t-}\subset\{(({\mathbf{X}}_{t^{\prime}},{\mathbf{Y}}_{t^{\prime}})\mid t\leq t-H\}$ as training samples and utilize the forecasting feedback (e.g., forecasting errors and gradients w.r.t. $\mathcal{D}_{t-}$) to update the model parameters. Their adapted parameters tend to align with the patterns or concepts present in ${\mathcal{D}}_{t-}$. However, it is crucial to recognize that the concepts present in ${\mathcal{D}}_{t-}$ may not necessarily reflect that of the test sample $({\mathbf{X}}_{t},{\mathbf{Y}}_{t})$ due the horizon time span $H$. To this end, the model optimized on ${\mathcal{D}}_{t-}$ might still be susceptible to concept drift, potentially resulting in suboptimal predictions at time $t$ (see details in Section 3).
@@ 3.1. Datasets
Following prior works (Zhang et al., 2023; Pham et al., 2023), we use five real-world time series datasets, including ETTh2, ETTm1, Weather, ECL, and Traffic. We provide their detailed descriptions in Appendix C.1 and statistical information in Table 7. We also adhere to the evaluation settings of FSNet (Pham et al., 2023) and OneNet (Zhang et al., 2023), where each dataset is chronologically divided into training/validation/test sets by the ratio of 20:5:75.
@@ 3.2. Baseline Variants
  • FSNet (Pham et al., 2023): FSNet is built upon a TCN (Bai et al., 2018) backbone and further develops an advanced updating structure that facilitates fast adaptation to concept drift.
  • OneNet (Zhang et al., 2023): OneNet is an ensemble model that dynamically combines two forecasters, one focused on temporal dependence and one focused on channel dependence. We follow its official implementation where the two forecasters are built upon FSNets.
  • PatchTST (Nie et al., 2023): PatchTST is one of the state-of-the-art time series forecasters, which models temporal patterns by Transformer.
@@ 3.3. Evaluation Metrics
Following previous works (Pham et al., 2023; Zhou et al., 2021; Zeng et al., 2023), we evaluate the forecasting performance by two commonly used metrics as defined below:
[…]
@@
  • We augment FSNet and OneNet with RevIN (Kim et al., 2022) for better performance. The results differ from those in the original paper. We explain the reasons in Appendix C.2.
@@ 4.1. Solution Overview
Following existing online model adaptation methods (Pham et al., 2023; Zhang et al., 2023), we use one training sample for model updating at each time $t$, i.e., $\mathcal{D}_{t-}={(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})}$. With the model updated on $\mathcal{D}_{t-}$, we estimate the concept drift between $\mathcal{D}_{t-}$ and $\mathbf{X}_{t}$, predict potential shifts in the parameter space, and accordingly make parameter adjustments. Formally, our Proceed solution consists of four key steps at each time $t$ as listed below.
@@ 4.2. Concept Drift Estimation
In the literature, there are numerous concept drift detection methods (Gama et al., 2014) that estimate the degree of concept drift to decide when to adapt the model. The degree can be estimated by the changes in forecast error (Caccia et al., 2020), distribution distance (Du et al., 2021), prediction uncertainty (Zhu et al., 2023), etc. Nevertheless, the degree as a scalar in $\mathbb{R}$ is not informative and cannot indicate how to adapt the model. Therefore, we propose to model a high-dimensional representation vector in $\mathbb{R}^{d_{c}}$ that characterizes both the degree and the direction of the concept drift, where $d_{c}$ is the representation dimension.
@@
Next, we estimate the concept drift between time $t-H$ and time $t$ by the concept difference, i.e., $\boldsymbol{\delta}_{t-H\to t}=\mathbf{c}_{t}-\mathbf{c}_{t-H}\in\mathbb{R}^{d_{c}}$.
@@ 4.3. Proactive Model Adaptation
(3) $[{\boldsymbol{\alpha}}_{t}^{(\ell)},{\boldsymbol{\beta}}_{t}^{(\ell)}]={\mathbf{W}_{2}^{(\ell)}}^{\top}\left(\sigma\left(\mathbf{W}_{1}^{(\ell)^{\top}}\boldsymbol{\delta}_{t-H\to t}+\mathbf{b}^{(\ell)}\right)\right)+\mathbf{1},$
@@ 4.4. Mini-batch Training
As our model adapter learns the transition probability $P(\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}_{s},\mathbf{X}_{s},\mathbf{X}_{t})$, we can follow the proof in the previous work (Bai et al., 2023) to show that it is feasible for $\widehat{\boldsymbol{\theta}}_{t}$ in Eq. (4) to yield lower forecasting error than $\boldsymbol{\theta}_{t-H}$ (see Appendix A).
@@
FSNet (Pham et al., 2023) $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})^{\dagger}$ GD + Adjusting convolutional filters based on $\nabla\text{MSE}(\hat{\mathbf{Y}}_{t-H},\mathbf{Y}_{t-H})$
OneNet (Zhang et al., 2023) $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})^{\dagger}$ GD + Adjusting ensemble weights based on $\text{MSE}(\hat{\mathbf{Y}}_{t-H},\mathbf{Y}_{t-H})$
SOLID (Chen et al., 2024) $\left\{(\mathbf{X}_{s},\mathbf{Y}_{s}):s\leq t-H\wedge\|\mathbf{X}_{s}-\mathbf{X}_{t}\|_{2}\in TopK\right\}$ GD
[…]

### v5

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3690624.3709210 $c$s: Information systems Data stream miningccs: Computing methodologies Online learning settings
@@ 1. Introduction
Time series forecasting has been a prevalent task in numerous fields such as climate (Pangu), energy (SolarNIPS; Sun2023DSE), retail (Boese2017), and finance (DoubleAdapt). Recent years have witnessed a surge of deep learning-based forecast models (PatchTST; iTransformer; GPT4TS; MOIRAI; SOFTS) that take past time series observations to predict values in the next $H$ steps, where $H$ is called forecast horizon. Due to the dynamic nature of the environment, latent concepts that influence observation values (e.g., social interest (DRAIN), stock market sentiment (DDGDA)) often change over time. This ubiquitous phenomenon is known as concept drift (SurveyConcepDrift). In the presence of concept drift, future test data may not follow a similar data distribution as the historical training data, causing degradation in forecasting performance (LLF; DoubleAdapt; duan2025brain).
@@
Online learning is commonly used to mitigate the effects of concept drift. The key consideration is to continually transform the newly observed time series into a set of training samples and adjust model parameters by minimizing the forecast errors on the new training samples. In addition to the standard fine-tuning technique, recent works (FSNet; OneNet) proposed more advanced model adaptation techniques, which focus on how to effectively adapt to recent data by utilizing forecasting feedback (e.g., errors or gradients) on the new training samples. Among them, FSNet (FSNet) monitors the gradients in previous fine-tuning, transforms them into parameter adjustments, and tailors a new forecast model to the current training samples. OneNet (OneNet) is an online ensembling network that generates ensemble weights to combine two forecast models and dynamically adjusts the ensemble weights and the models’ parameter weights according to the forecast errors. However, existing online learning methods overlook the fact that the ground truth of each prediction is not available immediately but is delayed after the forecast horizon.
@@
As illustrated in Fig. 1, an online learning task inherently has a $H$-step feedback delay in time series forecasting, resulting in a temporal gap (at least $H$ steps) between available training samples and the test sample. Formally, at each online time $t$, the current horizon window $\mathbf{Y}_{t}$, which has not been observed yet, has an overlap with $\mathbf{Y}_{t-H+1},\dots,\mathbf{Y}_{t-1}$. These samples cannot provide complete forecasting feedback and supervision for the performant, prevailing time series forecasting models that predict $H$-step values directly (DLinear; PatchTST; iTransformer). Hence, the available training samples for online learning at time $t$ is $\mathcal{D}_{t-}=\{(\mathbf{X}_{t^{\prime}},\mathbf{Y}_{t^{\prime}})\mid t^{\prime}\leq t-H\}$, which has a temporal distance to the test sample $(\mathbf{X}_{t},\mathbf{Y}_{t})$. The temporal gap issue suggests that addressing concept drift by fitting the forecast model to the recent training samples may be insufficient, since concept drift may also occur over the $H$-step temporal gap between the recent training samples and the current test sample.
@@
To investigate the impact of this temporal gap on forecasting performance, we examine two online learning strategies using real-world time series datasets (detailed in Section 3). The first strategy fine-tunes model parameters using the latest available training sample at time $t$, which is $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$. The second strategy (FSNet; OneNet) omits the feedback delay and adapts the model using $(\mathbf{X}_{t-1},\mathbf{Y}_{t-1})$, which is infeasible in practice due to future information leakage. The empirical results show that the average forecast error of the first strategy is approximately double that of the second strategy, indicating that even the latest observed time series pattern differs from the test sample notably. Moreover, the performance gap between the two strategies becomes more significant with a longer horizon, suggesting that the concept drift becomes more pronounced over longer time intervals. These findings highlight the presence of a substantial concept drift between the practical training samples and the test sample, limiting the effectiveness of current online learning techniques that adapt model parameters to potentially outdated concepts and leave the concept drift unresolved.
@@
In this paper, we aim to answer the question: how can we effectively adapt the forecast model to each test sample and boost online time series forecasting performance against ongoing concept drifts? A straightforward idea is to estimate a sketch of the latent concept of the test data and customize model parameters for the estimated concept. However, it is challenging to train a model adapter (e.g., a mapping function) that directly generates model parameters based on the estimated concept of the test sample. First, test data may bear a new concept that is out of the distribution of the historical time series. There is no experience in generating optimal parameters for such an out-of-distribution concept. Second, a parameterized model adapter that maps concepts to model parameters (or their updates) essentially lies in a parameter space of $\mathbb{R}^{m\times n}$, where $m$ is the concept embedding dimension and $n$ is the number of model parameters. Since advanced time series forecasting models often involve a high quantity of parameters (e.g., 1 million (PatchTST; iTransformer)), the model adapter is hard to optimize and easily suffers from overfitting due to the huge parameter size and limited test data.
To address these issues, we propose Proceed, a PROaCtivE modEl aDaptation framework that responds to the concept drift before forecasting the test sample. In scenarios where optimal model parameters change with time, we posit that such parameter changes are affected by the drift across a latent concept space, i.e., there may exist learnable relationships between the direction and degree of concept drift and those of parameter changes. In light of this, we propose to infer parameter changes based on concept drift, rather than directly generating a whole model. Specifically, Proceed begins with a forecast model that has learned new training samples. Given a test sample, Proceed exploits latent features that encapsulate the concept drift between the training samples and the test sample. In response to the undergoing concept drift, Proceed generates parameter adjustments via bottleneck layers, which yield a compact set of adaptation coefficients that rescale all model parameters. In this way, our solution eschews direct mapping from concept drift to per-parameter adjustments, instead opting for a more nuanced and efficient strategy. We expect the adapted model with rescaled parameters to approach the optimum for the test sample. To enhance the generalization ability of Proceed, we shuffle historical data to synthesize diverse concept drifts, on which we train Proceed to learn the relationships between concept drifts and desirable parameter adjustments. When confronted with a reoccurring concept drift that has been learned among the synthetic concept drifts, Proceed can tailor the model to the test sample.
@@
  • We propose Proceed, a proactive model adaptation framework for online time series forecasting under concept drift. Proceed first estimates the latent representation of concept drift and then customizes parameter adjustments that adapt the parameters to the concept of the test sample in advance of online prediction.
@@ 2. Definitions and Problem
Typically, in multi-step forecasting where $H>1$, there are two primary strategies to generate the predictions ${\bf\hat{Y}}_{t}$ at each time $t$. The first strategy performs iterative forecasting. That is, the model forecasts one step ahead and uses its prediction to forecast the next step, repeating this process for the entire horizon. Despite its simplicity, this strategy suffers from significant error accumulation over long horizons (DLinear). The second strategy adopts direct forecasting where the model makes all the predictions ${\bf\hat{Y}}_{t}$ simultaneously. Note that the two strategies require different output modules or layers in the forecast model. Recent works (DLinear; PatchTST) have shown that direct forecasting tends to outperform the iterative method, particularly for longer forecast horizons. Hence, this paper adopts the direct forecasting strategy where a sample $\mathbf{X}_{t}$ is considered valid for training if all the $H$-step values in $\mathbf{Y}_{t}$ are known (SOLID).
In online forecasting scenarios, time series data are observed sequentially. Due to the dynamic nature of real-world processes, underlying data distributions are subject to constant change. Consequently, a forecast model trained on historical data may encounter difficulties when confronted with new, evolving patterns. This issue is referred to as the concept drift challenge, which can substantially affect model performance over time (SurveyConcepDrift). To mitigate concept drifts, it is crucial to adapt the forecast model continuously to assimilate new concepts presented in the incoming time series. Formally, the online model adaptation problem is defined as follows.
@@
At a high level, existing model adaptation methods (FSNet; OneNet; SOLID) select some observed data $\mathcal{D}_{t-}\subset\{(({\mathbf{X}}_{t^{\prime}},{\mathbf{Y}}_{t^{\prime}})\mid t\leq t-H\}$ as training samples and utilize the forecasting feedback (e.g., forecasting errors and gradients w.r.t. $\mathcal{D}_{t-}$) to update the model parameters. Their adapted parameters tend to align with the patterns or concepts present in ${\mathcal{D}}_{t-}$. However, it is crucial to recognize that the concepts present in ${\mathcal{D}}_{t-}$ may not necessarily reflect that of the test sample $({\mathbf{X}}_{t},{\mathbf{Y}}_{t})$ due the horizon time span $H$. To this end, the model optimized on ${\mathcal{D}}_{t-}$ might still be susceptible to concept drift, potentially resulting in suboptimal predictions at time $t$ (see details in Section 3).
@@ 3.1. Datasets
Following prior works (OneNet; FSNet), we use five real-world time series datasets, including ETTh2, ETTm1, Weather, ECL, and Traffic. We provide their detailed descriptions in Appendix C.1 and statistical information in Table 7. We also adhere to the evaluation settings of FSNet (FSNet) and OneNet (OneNet), where each dataset is chronologically divided into training/validation/test sets by the ratio of 20:5:75.
@@ 3.2. Baseline Variants
  • FSNet (FSNet): FSNet is built upon a TCN (TCN) backbone and further develops an advanced updating structure that facilitates fast adaptation to concept drift.
  • OneNet (OneNet): OneNet is an ensemble model that dynamically combines two forecasters, one focused on temporal dependence and one focused on channel dependence. We follow its official implementation where the two forecasters are built upon FSNets.
  • PatchTST (PatchTST): PatchTST is one of the state-of-the-art time series forecasters, which models temporal patterns by Transformer.
@@ 3.3. Evaluation Metrics
Following previous works (FSNet; Informer; DLinear), we evaluate the forecasting performance by two commonly used metrics as defined below:
$\text{Mean Square Error (MSE)}=\frac{1}{N|T_{\text{test}}|}\sum_{t\in T_{\text{test}}}\|\hat{\mathbf{Y}}_{t}-{\mathbf{Y}}_{t}\|_{2}^{2},$
$\text{Mean Absolute Error (MAE)}=\frac{1}{N|T_{\text{test}}|}\sum_{t\in T_{\text{test}}}\|\hat{\mathbf{Y}}_{t}-{\mathbf{Y}}_{t}\|_{1},$
@@
  • We augment FSNet and OneNet with RevIN (RevIN) for better performance. The results differ from those in the original paper. We explain the reasons in Appendix C.2.
@@ 4.1. Solution Overview
Following existing online model adaptation methods (FSNet; OneNet), we use one training sample for model updating at each time $t$, i.e., $\mathcal{D}_{t-}={(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})}$. With the model updated on $\mathcal{D}_{t-}$, we estimate the concept drift between $\mathcal{D}_{t-}$ and $\mathbf{X}_{t}$, predict potential shifts in the parameter space, and accordingly make parameter adjustments. Formally, our Proceed solution consists of four key steps at each time $t$ as listed below.
@@ 4.2. Concept Drift Estimation
In the literature, there are numerous concept drift detection methods (SurveyConcepDrift) that estimate the degree of concept drift to decide when to adapt the model. The degree can be estimated by the changes in forecast error (OSAKA), distribution distance (AdaRNN), prediction uncertainty (METER), etc. Nevertheless, the degree as a scalar in $\mathbb{R}$ is not informative and cannot indicate how to adapt the model. Therefore, we propose to model a high-dimensional representation vector in $\mathbb{R}^{d_{c}}$ that characterizes both the degree and the direction of the concept drift, where $d_{c}$ is the representation dimension.
@@
Next, we estimate the concept drift between time tHt-H and time $t$ by the concept difference, i.e., $\boldsymbol{\delta}_{t-H\to t}=\mathbf{c}_{t}-\mathbf{c}_{t-H}\in\mathbb{R}^{d_{c}}$.
@@ 4.3. Proactive Model Adaptation
(3) ${\boldsymbol{\alpha}}_{t}^{(\ell)},{\boldsymbol{\beta}}_{t}^{(\ell)}]={\mathbf{W}_{2}^{(\ell)}}^{\top}\left(\sigma\left(\mathbf{W}_{1}^{(\ell)^{\top}}\boldsymbol{\delta}_{t-H\to t}+\mathbf{b}^{(\ell)}\right)\right)+\mathbf{1},$
@@ 4.4. Mini-batch Training
As our model adapter learns the transition probability $P(\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}_{s},\mathbf{X}_{s},\mathbf{X}_{t})$, we can follow the proof in the previous work (DRAIN) to show that it is feasible for $\widehat{\boldsymbol{\theta}}_{t}$ in Eq. (4) to yield lower forecasting error than $\boldsymbol{\theta}_{t-H}$ (see Appendix A).
@@
FSNet (FSNet) $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})^{\dagger}$ GD + Adjusting convolutional filters based on $\nabla\text{MSE}(\hat{\mathbf{Y}}_{t-H},\mathbf{Y}_{t-H})$
OneNet (OneNet) $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})^{\dagger}$ GD + Adjusting ensemble weights based on $\text{MSE}(\hat{\mathbf{Y}}_{t-H},\mathbf{Y}_{t-H})$
SOLID (SOLID) $\left\{(\mathbf{X}_{s},\mathbf{Y}_{s}):s\leq t-H\wedge\|\mathbf{X}_{s}-\mathbf{X}_{t}\|_{2}\in TopK\right\}$ GD
Proceed $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$ GD Adjusting all parameters based on $\mathbf{X}_{t-H}$ and $\mathbf{X}_{t}$